OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of -5 and a y-intercept of -1.

OpenStudy (anonymous):

y=-5/1x-1

OpenStudy (anonymous):

the slope is always a fraction,

OpenStudy (anonymous):

oh ok could you kinda show me how you got it ?

OpenStudy (anonymous):

so, the slope always has an x next to it. Slope intercept formula is y=mx+b right?

OpenStudy (anonymous):

the m is the slope you are given, and the b is the y-intercept, or where the line crosses the y axis

OpenStudy (anonymous):

oh i get it thank you can you help with one more please?

OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 3x - 3y = 4. ?

OpenStudy (anonymous):

sure! so the first thing you have to do is arrange the equasion into slope intercept form

OpenStudy (anonymous):

so arrange the given question into y=mx+b

OpenStudy (anonymous):

i dont really know what they each stand for...

OpenStudy (anonymous):

im taking the pretest now

OpenStudy (anonymous):

lol m stands for slope, and b stands for y intercept

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

anything else?

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. just tell me if im rigt is it

OpenStudy (anonymous):

y = x + 2?